Calculate the enthalpy of the methanol combustion reaction using the binding energies.


2CH3OH (I) + 3O2 (g) → 2CO2 (g) + 4H2O (I)

1
Expert's answer
2021-02-01T04:04:28-0500

Solution:

2CH3OH(I) + 3O2(g) → 2CO2(g) + 4H2O(I)

To find enthalpy of the methanol combustion reaction, use the following formula:

ΔHrxn = ∑(bonds broken) − ∑(bonds formed)


In this reaction, 6 C-H bonds, 2 C-O bonds, 2 O-H bonds and 3 O=O bonds must be broken.

Also, 4 C=O bonds and 8 O-H bonds are formed.


Bond energies of these bonds:

  • C-H bond: 413 kJ/mol
  • C-O bond: 358 kJ/mol
  • O-H bond: 463 kJ/mol
  • O=O bond: 495 kJ/mol
  • C=O bond in CO2: 799 kJ/mol


Bonds broken = 6(C-H) + 2(C-O) + 2(O-H) + 3(O=O)

Bonds broken = 6×(413 kJ/mol) + 2×(358 kJ/mol) + 2×(463 kJ/mol) + 3×(495 kJ/mol) = 5605 kJ/mol

Bonds broken = 5605 kJ/mol


Bonds formed = 4(C=O) + 8(O-H)

Bonds formed = 4×(799 kJ/mol) + 8×(463 kJ/mol) = 6900 kJ/mol

Bonds formed = 6900 kJ/mol


ΔHrxn = Bonds broken - Bonds formed = 5605 kJ/mol - 6900 kJ/mol = -1295 kJ/mol

ΔHrxn = -1295 kJ/mol


Answer: The enthalpy of the methanol combustion reaction (ΔHrxn) is -1295 kJ/mol.